OpenStudy (ragingsquirrel):

What is the area of the trapezoid? 88 in2 128 in2 96 in2 48 in2

OpenStudy (ragingsquirrel):



Directrix (directrix):

A = (1/2) h ( sum of bases)

Directrix (directrix):

A = (1/2) (8) (6+16) A = 88 square units

Directrix (directrix):

Questions?

OpenStudy (ragingsquirrel):

>.> I'll just take a guess here...

OpenStudy (pokemon23):

like directix

OpenStudy (pokemon23):

We used the trapezoid formula 1/2h(b1+b2)

OpenStudy (pokemon23):

In this case the height is 8 and the bases are 6 and 16